OpenStudy (anonymous):

There are values of t so that the equation cos t = 3/5 holds.

OpenStudy (anonymous):

@beccaboo333 can help you

OpenStudy (anonymous):

true or false?

OpenStudy (anonymous):

i think false

OpenStudy (anonymous):

:-):-(

OpenStudy (anonymous):

@sourwing do you know

OpenStudy (anonymous):

correct \[\cos t=\frac{ 3 }{ 5 }<\left| 1 \right|\] it is true.